Water discharge through a pipe of length 50m and diameter 0.5m at 3m3 /s. For this flow, f = 0.025 . Determine the head loss through the model pipe for a scale ratio of 1:10.


1
Expert's answer
2021-08-16T02:32:56-0400

"h_L = (f\dfrac{L}{d}+ \Sigma K_L)\dfrac{v²}{2g}"


"v = \dfrac{V}{A} =\dfrac{3}{π(0.5²)/4}= 15.28m/s"


"h_L = (0.025(\dfrac{50}{0.5}) +\dfrac{1}{10}(50))\dfrac{15.28²}{2×9.8}"

"=89.34m"


"\therefore" The head loss through the pipe is 89.34m